OpenStudy (anonymous):

Each side of triangle A is 10 inches long. Each side of triangle B is 15 inches long. The area of triangle A is about 43.3 square inches. What is the approximate area of triangle B?

OpenStudy (anonymous):

10/15 = 43.3/x 10/15 = 43.3/64.95 64.95 inches is the answer

OpenStudy (anonymous):

That is how I set it up as well, but the answer was 97.4 inches ^2.

OpenStudy (anonymous):

triangle B would be 64.95 because u do 10/43.3 = 15/x Then cross multiply and divide (10x = 43.3(15))

OpenStudy (anonymous):

I think you have to convert the side length into something for inches square because the problem is asking for the area in inches square, and the sides are in inches, so I don't think you can use cross multiplication.